OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope. (7, –4); m = 5/3 my answer. y-7=5/3(x+4) ????

OpenStudy (amistre64):

your 7 and 4 are misplaced

OpenStudy (amistre64):

remember P(x=7, y=-4)

OpenStudy (amistre64):

y-Py = m(x-Px) y-(-4) = (5/3)(x-(7)) \[y+4 = \frac{5}{3}(x-7)\]